OpenStudy (anonymous):

Please Help!! Multiply (-2n)^2 times (3n^6)^2=?

OpenStudy (anonymous):

\((-2n)^2=4n^2\) and \(( 3n^6)^3=9n^{12}\) for a start

OpenStudy (anonymous):

ok typo on the last one should be \[(3n^6)^2=9n^{12}\]

OpenStudy (anonymous):

in any case your job is to multiply \[4n^2\times 9n^{12}\] multiply \(4\times 9\) and add the exponents to get your answer

OpenStudy (anonymous):

ok